OpenStudy (anonymous):

Solve the system of equations using substitution. A) (2,8) B) (2,-8) C) (10,24) D) no solution

OpenStudy (anonymous):

where's the graph or is their none

OpenStudy (anonymous):

@Esi7023

OpenStudy (texaschic101):

equations please

OpenStudy (camerondoherty):

What system of equations O.o

OpenStudy (anonymous):

not 100%sur but I think D but hold on

OpenStudy (texaschic101):

ok...you see y = 2x + 4 so we then sub 2x + 4 in for y in the other equation 2x - y = 6 2x - (2x + 4) = 4 -- distribute through the parenthesis 2x - 2x - 4 = 4 --- add 4 to both sides 2x - 2x = 4 + 4 0 = 8 no solution

OpenStudy (anonymous):

so yea D

OpenStudy (texaschic101):

yep..I agree with you :)

OpenStudy (anonymous):

your answer is D @Esi7023

OpenStudy (anonymous):

thxs:)